OpenStudy (anonymous):

-4 (+-) sqrt -44 -------------- 6 Whats next?

OpenStudy (anonymous):

How did you get?\[\sqrt{-44}\]You can't square root a negative, check your math again.

OpenStudy (anonymous):

3x2 + 4x = -5

OpenStudy (anonymous):

came from this using the quadratic equation

OpenStudy (anonymous):

\[x=\frac{-4 \pm \sqrt{-44}}{6}\]\[x=\frac{-4 \pm 2\sqrt{-11}}{6}\]\[x=\frac{-2 \pm i\sqrt{11}}{3}\]

OpenStudy (anonymous):

x =\[(-4\pm \sqrt{-44})/6\]

OpenStudy (anonymous):

\[\sqrt{-1}=i\] This answer is in the complex plane

OpenStudy (anonymous):

\[3x^2+4x-5=0\]Quadratic Equation\[((-4)+-\sqrt{(4)^2-4(3)(5)})/2(3)\]\[(-4+-\sqrt{-44})/6\]Zed's right, complex plane. I hate imaginary numbers.

OpenStudy (anonymous):

Zed, how do you get your equations all over 6 like that?

OpenStudy (anonymous):

\[(-4\pm \sqrt{44(i ^{2})})/6\] \[(-4\pm i \sqrt{44})/6\] \[(-4\pm i \sqrt{2*2*11})/6\] \[(-4\pm 2i \sqrt{11})/6\] Take 2 as common from the numerator \[2(-2\pm i \sqrt{11})/6\] \[(-2\pm i \sqrt{11}) / 3\]